Suppose that while a mining cart is rolling along a straight, frictionless, horizontal track with some initial speed, it starts to rain. If the rain falls vertically into the cart and begins to accumulate within it, flooding the cart, determine the following

19. is the momentum of the system conserved in the horizontal direction?

a. Yes

b. No

20. Is the momentum of the system conserved in the vertical direction?

a. Yes

b. No

22.If a 1200 kg train car is initially moving with a speed of 40 m/s, calculate the force needed to bring this car to rest in a distance of 200 meters and a time of 10 s.

a. 6000 N

b. 240 kN

c. 4800 N

d. 120N

23. Find the torque exert about a bolt by a wrench when a 90 N force is exerted at an angle of 80° to the wrench at a distance of 15 cm from the bolt's center.

a. 1350 Nm

b. 13.5 Nm

c. 13.3 Nm.

d. 1330 Nm.
Answer #1

here,

19)

a)

yes momentum will bw conserved in horizontal direction

20)

a)

yes momentum will bw conserved in vertical direction

21)

m1 = 1200 kg

speed , v = 40 m/s

s = 200 m and t = 10 s

let the accelration be a

0 = 40 + a * 10

a = - 4 m/s^2

the force needed , F = m * a = c) 4800 N

22)

force , F = 90 N

r = 0.15 m

theta = 80 degree

the torque , t = F * r * sin(80)

t = 90 * 0.15 * sin(80)

t = c) 13.3 N.m